x+2y=-1
2x-3y=12
Solve using substitution.

Answer by stanbon(75887)

x+2y=-1
2x-3y=12
Solve using substitution.
-----
x = -2y-1
----
Substitute for "x" and solve for "y":
2(-2y-1) - 3y = 12
----
-4y -2 -3y = 12
-----
-7y = 14
y = -2
----
Solve for "x":
x = -2y-1
x = -2*-2-1
x = 3
--------------------
Cheers,
Stan H.
====================